OpenStudy (kpop4life123):

A waffle cone has a height of 7 inches and a diameter of 3 inches. What is the volume of ice cream that can be contained within the cone? Use 3.14 for pi. Round your answer to the nearest hundredth.

OpenStudy (kpop4life123):

20.45 cubic inches 21.34 cubic inches 22.44 cubic inches 23.78 cubic inches

OpenStudy (anonymous):

hey buddy haha need so me help?

OpenStudy (chealyn98):

Volume of a cone is given by 1/3 * pi * r * r * h r = d/2 r = radius d = diameter h = height

OpenStudy (kpop4life123):

i dont know it :/

OpenStudy (anonymous):

as chealyn said, you need to find the volume of the cone with the equation..

OpenStudy (kpop4life123):

how lol im really bad at math

OpenStudy (chealyn98):

I think it's 1/3 because there's three sides in a triangle and a cone is just a 3d form of a triangle. The formula is what gave you so all you have to do is plug in the numbers in your calculator (or however you get your answers) and you'll get it. To find the radius when given the diameter is simple (just divide by two) Explained Answer: 1/3 * 3.14 * 1.5 * 1.5 * 7 = 16.485 Which can be rounded to --> 16.49 <--

OpenStudy (anonymous):

That's correct @Chealyn98

OpenStudy (kpop4life123):

THANK YOU FOR SHOWING ME HOW TO DO IT AND FOR THE ANSWER!!!!!!!!!!

OpenStudy (chealyn98):

Your welcome

OpenStudy (chealyn98):

If you need any more help just let me know. 👍